# Nonlinear phase synchronization and the role of spacing in shell models

### v1

@@ Abstract
A shell model can be considered as a chain of triads, where each triad can be interpreted as a nonlinear oscillator that can be mapped to a spinning top. Investigating the relation between phase dynamics and intermittency in a such a chain of nonlinear oscillators, it is found that synchronization is linked to increased energy transfer. In particular, the results provide evidence that the observed systematic increase of intermittency, as the shell spacing is decreased, is associated with strong phase alignment among consecutive triadic phases, facilitating the energy cascade. It is shown that while the overall level of synchronization can be quantified using a Kuramoto order parameter for the global phase coherence in the inertial range, a local, weighted Kuramoto parameter can be used for the detection of burst-like events propagating across shells in the inertial range. This novel analysis reveals how partially phase-locked states are associated with the passage of extreme events of energy flux. Applying this method to helical shell models, reveals that for a particular class of helical interactions, a reduction in phase coherence correlates with suppression of intermittency. When inverse cascade scenarios are considered using two different shell models including a non local helical shell model, and a local standard shell model with a modified conservation law, it was shown that a particular phase organization is needed in order to sustain the inverse energy cascade. It was also observed that the PDFs of the triadic phases were peaked in accordance with the basic considerations of the form of the flux, which suggests that a triadic phase of $\pi/2$ and $-\pi/2$ maximizes the forward and the inverse energy cascades respectively.
@@ I Introduction
Many natural systems exhibit non-Gaussian fluctuations across a broad range of spatial and temporal scales. This implies the presence of extreme fluctuations or rare events, whose dynamical properties and statistical signatures become essential to describe the overall behavior of the system. Examples include seismic activity, financial market crashes and extreme weather events [1, 2, 3]. An archetypical example of such complex, open system is the problem of fully developed turbulence: an out of equilibrium state of fluid or plasma flow, displaying a mixture of hierarchical self organization and chaotic behavior across a large range of scales, maintained in a statistically stationary state by continuous energy injection (e.g. at large scales) and dissipation (e.g. at small scales).
@@
Between forcing and dissipation, turbulent systems can develop an inertial range, spanning many decades, exhibiting self-similar properties heuristically captured by the Richardson cascade picture, and the power law spectrum as predicted by the Kolmogorov theory [4].
@@
Since direct numerical simulations (DNS) of turbulent flows over a wide range of scales were historically too costly, various reduced models were developed in the past [5] that are capable of reproducing various statistical features of the full system such as the power law distribution of energy as a function of scale and the scaling of higher-order structure functions. Among those, shell models[6, 7] stand out because of the versatility in their ability to mimic certain statistical aspects of different types of turbulent systems including rotating turbulence, passive scalar advection, thermal convection etc. [8, 9, 10]. In the context of plasma physics, shell models have been used to describe magnetohydrodynamic (MHD) turbulence [11, 12], and drift waves in fusion plasmas [13, 14].
@@
Probably the most intriguing – perhaps unexpected – feature of shell models, is their ability to reproduce the anomalous scaling of the structure functions similar to fully developed turbulence [15, 16, 17]. This makes shell models a useful tool to capture the statistics of extreme events [18, 19], which are manifested as burst like structures, associated to finite time singularities or instantons [20, 21], playing an important role in shaping the underlying intermittency characteristics.
@@
Shell models reduce the complexity of the nonlinear energy transfer governed by triadic interactions drastically, while preserving the key properties of the underlying system such as its scale invariance, the quadratic nature of the nonlinearity and the conservation laws - like Energy and Helicity in the case of three dimensional turbulence. This extreme reduction in the number of degrees of freedom, is achieved by dividing the Fourier space into a set of geometrically spaced shells and retaining the nonlinear interactions that involve only consecutive shells, thus replacing the turbulent system with a one dimensional chain of connected, self-similar triads.
@@
In order to reintroduce some spatial structure in cascade models, various more sophisticated constructions have been proposed, including nested polyhedra [22], logarithmic discretization LDM [23], spiral chains [24] and logarithmic lattices [25, 26, 27]. Most of these models are self-similar by construction, and maintain a constant number of triads across different scales. However, a puzzling feature of these spatially-extended “shell” models is the apparent suppression of intermittency observed in their statistics. Similar loss of intermittency has also been observed in numerical simulations of fractally-decimated 3D turbulence [28] even keeping an overwhelming majority of the modes and dropping only a small percentage.
@@
The effect of a fractal decimation of Fourier modes has been explored also in simpler systems, such as the Burgers equation, where a disruption of the average phase coherence is correlated with a reduction of the intermittency, normally associated with shock formation [29, 30]. This suggests a picture, in which phase alignment between triads enhances energy transfer resulting in strong dissipation events [31, 32]. Generalizing this requires understanding the self-organization of triadic phases in fully developed turbulence, and the definition of a clear order parameter that identifies the interactions involved in the collective behavior responsible for intermittency, which remains a challenge. Part of the difficulty involves dealing with the huge number of degrees of freedom represented by the convolution sum in such a system. One possible way forward may be to reformulate turbulence as a network of Fourier modes [33, 34, 35], which allows us to consider reduced models, such as shell models as reductions of a complete grid into a network of groups of nodes, in this case represented by shells. While the actual complex phases of a shell model are not physical in any way, triadic phase coherences among its elements would still suggest some level of triadic phase coherence among the underlying Fourier modes that are lumped together to form the shells.
@@
Thus, in order to shed light on the link between intermittency and phase coherency in fully developed turbulence, we investigate this connection in shell models that preserve the same invariants as the Navier-Stokes system, and present, in some cases, a level intermittency comparable to that of 3D turbulence. In order to understand this, we use our earlier observation that decreasing inter-shell spacing in standard shell models increases intermittency considerably [36], which is also in accordance with the observations in studies of the Burger’s equation cited above.
@@
Notice that, a shell model can be interpreted as a one dimensional chain of connected triads, each representing a nonlinear oscillator. Consecutive triads only differ by a scaling factor related to inter-shell spacing. This means that by decreasing the inter-shell spacing, (while keeping the conserved quantities unchanged) both the number and the similarity of the oscillators would be increased making it potentially easier to form a phase coherent state. To quantify global phase coherence, and its relation to different levels of anomalous scalings of the structure functions, we employ a global Kuramoto order parameter, which allows us to show that higher intermittency correlates with enhanced phase coherence.
@@
In addition, we observe that phase coherence correlates with the existence of coherent structures, which take the form of localized pulses in shell models. In order to clarify the role of these localized events, we introduce a novel, local order parameter, in the form of a weighted Kuramoto parameter, which allow us to observe burst-like events propagating across shells in the inertial range and analyze the alignment of triadic phases in forming an “open channel” for energy transfer during these events. We show that extreme energy bursts are consistently preceded by rapid growth in local phase coherence and that more intermittent models display stronger and longer-lasting phase-locked states.
@@
Finally, we consider inverse cascades in shell models, which -apart from particular cases- are usually suppressed by equipartition solutions. We consider these special cases that actually work, in order to study the role of phase coherence in inverse cascade and to compare it with the forward cascade. The results reveal that, in the forward and inverse cascade regimes, the PDFs of the triadic phases exhibit peaks at values of $\pi/2$ and $-\pi/2$ respectively, which maximize the flux term at fixed amplitude. Notably, in the case of intermittent cascades the peak at $\pi/2$ , is much more marked, indicating that coherent events are clearly favored by the system; in contrast, for inverse cascades, where the distribution is symmetric.
@@
In section II we introduce the shell model framework, focusing on the dynamical equations for phases and amplitudes II.1, discussing role of phase alignments in the energy flux II.2 and presenting the phase statistics for the standard forward intermittent cascade. Next in II.3, we introduce the Kuramoto order parameters to quantify phase coherence in shell models, highlighting how local synchronization (sync) is related to extreme flux events.
@@
In Sec. IV, we extend the analysis to helical shell models, exploring how synchronization properties vary with shell spacing, focusing on a helical shell model exhibiting suppression of intermittency. Section V is devoted to role of phase organization in driving inverse cascades, we present an inverse energy cascade for an elongated helical shell model analyze the inverse energy cascade in a local shell model.
@@
II Shell model as a dynamical equations for phases and amplitudes
Motivated by observations of statistical isotropy and scale invariance of the turbulent cascade of fully developed turbulence without obvious sources of anisotropy, the shell model reduction consists in modeling the velocity field as a one-dimensional set of complex variables representing different scales, corresponding to a set of logaritmically spaced wave numbers $k_{n}=k_{0}g^{n}$, where it is customary to take the spacing factor $g=2$.
The evolution equation for the shell variable $u_{n}$, for the GOY (Geldzer-Okhitani-Yamada) model, which is the standard example for shell models, is usually written as
@@ II Shell model as a dynamical equations for phases and amplitudes
where the first line on the right hand side represents the reduction of the convolution sum to a local triadic interaction between three consecutive shells, $d_{n}=\nu k_{n}^{2p}+\mu k_{n}^{-2q}$ is the general form of (hyper/hypo)-viscosity, and $f_{n}$ correspond to an external forcing localized at a shell $n_{f},$or a small set of consecutive shells around it, corresponding to the typical energy injection scale $1/k_{n_{f}}$.
Here the coefficients $a,b,c$ are chosen in such a way that they preserve the typical inviscid invariants of the underlying physical system. In the case of 3D Navier-Stokes turbulence, it is the energy and the helicity that are conserved, which are defined for the shell models respectively as
@@
Enforcing the conservation of these two quantities one obtains, $a=g+1$, $b=g^{-1}-g$ and $c=-1-g$, as a function of the inter shell spacing $g$. It is also common to factor out $ak_{n+1}$ in the nonlinear term of Eq. (1) so that the other two coefficients corresponding to $b$ and $c$ can be written as $\left(g^{-2}-g^{-1}\right)$ and $-g^{-3}$ respectively. The common choice of $g=2$ would then give $-1/4$ and $-1/8$ for the second and third coefficients respectively. Another common choice in the literature is $g=\varphi$ ( the golden ratio), which gains significance, from the derivation of shell models based on a decimation of Fourier space, which requires the existence of triadic interations between the characteristic wave numbers of the shells (specifically solution to $\boldsymbol{k}_{n-1}+\boldsymbol{k}_{n}+\boldsymbol{k}_{n+1}=0$, which imposes the geometrical constrain $g\leq\varphi$ ). While there are some examples of studies of shell models varying their nonlinear coupling coefficients [38, 39], most of these studies are done in such a way that varying the coefficients change the conservation law, and hence the effective dimensionality of the underlying physical model, instead of keeping the conservation laws fixed and varying the logarithmic spacing for a fixed model. Exploring different shell spacings is also important for making the connection to a perspective based on Fourier space decimation, such as logarithmic lattices [26], LDM [23], or helical shell models with different types of elongations [40].
@@
Note that a standard shell model with $N$ nodes has $N-2$ interacting triads. Since the equation for a single unconnected triad is equivalent to that of a spinning top, giving rise to nonlinear oscillations that can be represented in terms of Jacobi elliptic functions, each triad can be considered as a nonlinear oscillator. Furthermore since the shell model consists of self similar triads that are scaled by the same factor, the consecutive nonlinear oscillators that are coupled are also self-similar as they are also scaled by the same factor. The flux through such a chain of triads then requires some sort of synchronization of the phases of those consecutive triads. Decreasing the shell spacing is thus key to understanding the synchronization behavior of shell models. Since it simultaneously allows us to increase $N$, considering the large $N$ limit, important for synchronization, while at the same time reducing the ratio of the coefficients of two consecutive triads allowing them to have closer natural nonlinear frequencies. Note that the usual GOY model with $g=2$, covers roughly seven decades in $k$-space with $N=24$, and it is not feasible to increase that number substantially without decreasing $g$, since the power law solution would start going under machine precision levels.
@@
Thus, varying the $g$ parameter, while increasing $N$ in order to cover the same wave-number range, allows us to modify the affinity of the triadic network of the shell model to synchronization (i.e. synchronizability). Since the intermittency in shell models is a consequence of its time evolution, one would expect this to have significant impact on intermittency as well. Based on these observations, we present below, a detailed study introducing specific methods for observing various aspects of local and global synchronization, and the implications of these observations for the turbulent cascade as well as intermittency, mainly by varying the $g$ parameter to control its sychronizability.
II.1 Basic Phase Dynamics in Shell Models
@@
that leaves triadic phases invariant for any arbitrary parameter $\delta.$ The consequence of this $U(1)$ symmetry is that single shell phases $\theta_{n}$ are not uniquely determined by the triadic phases, and as also numerically confirmed, they are uniformly distributed. In other words, the appropriate quantities, whose dynamics and statistics are physically relevant are the triadic phases $\varphi_{n}$, rather than single shell phases $\theta_{n}$.
@@
$+a\Lambda_{n+2}^{n+1}\cos\varphi_{n+2}+b\Lambda_{n+1}^{n+1}\cos\varphi_{n+1}+c\Lambda_{n}^{n+1}\cos\varphi_{n}.$
@@
$\Lambda_{\ell}^{n}\equiv k_{\ell}\frac{U_{\ell-1}U_{\ell}U_{\ell+1}}{U_{n}^{2}}$
Note that Eq. (5) involves the triadic phases of the five consecutive triads from $n-2$ to $n+2$, which in turn include contributions from seven consecutive shells from $n+3$ to $n-3$, making this coupling longer range compared to the shell amplitude equation, which involves only the neighboring triadic phases (i.e. $\varphi_{j}$ with $j=n-1,n,n-1$), including contributions from shells from $n-2$ to $n+2$. Note that these couplings can also be interpreted (and generalized to more complex interactions) within the network representation of the shell models, see e.g [34], where the evolution of a given triad can be computed as a sum over all the triads connected through shared nodes in the network topology.
II.2 Phase and Flux
The phase dynamics of a shell model with $N$ shells can therefore be considered as being equivalent to that of a chain of $N-2$ nonlinear oscillators representing the triads, where the coupling terms are determined by a combination of the $N$ self-consistently evolving amplitudes and the interaction coefficients. When we introduce forcing near the first few shells and dissipation at high $n$, the chain of oscillators of the shell model are constrained to accommodate the flow of energy from low to high $n$. If the phases are initially distributed in such a way that there is no flux, the amplitudes will grow up to the point that the phases will organize themselves in order to accommodate the spectral flux implied by forcing and dissipation, basically because of the free energy implied in such a non-uniform distribution of energy. This means that the energy injection at low $n$ will drive phases to dynamically self-organize in order to sustain the cascade mechanism, possibly resulting in an eventually constant energy flux.
In order to see this steady state, let us consider the form of the flux in a shell model. The energy cascades to small scales due to nonlinear interactions, with the energy transfer in a given triad centered at shell $n$, defined as a local spectral energy flux, given in terms of the three shell correlation $\chi_{n}=U_{n-1}U_{n}U_{n+1}e^{-i\varphi_{n}}$ as:
$\begin{array}[]{cc}\Pi_{N}^{E}=&2\text{Im}\left(ak_{n+1}\chi_{n+1}-ck_{n}\chi_{n}\right)\end{array}$ (6)

### v2

@@ Abstract
A shell model can be considered as a self-similar chain of interacting triads, where each triad can be interpreted as a nonlinear oscillator that can be mapped to a spinning top. Investigating the relation between phase dynamics and intermittency in such a chain of nonlinear oscillators, it is found that synchronization is linked to increased energy transfer. In particular, our results indicate that the observed systematic increase of intermittency, as the shell spacing is decreased, is associated with strong phase alignment among consecutive triadic phases, facilitating the energy cascade. It is shown that while the overall level of synchronization can be quantified using a Kuramoto order parameter for the global phase coherence in the inertial range, a local, weighted Kuramoto parameter can be used for the detection of burst-like events propagating across shells in the inertial range. This novel analysis reveals how locally phase-locked states are associated with the passage of extreme events of energy flux. Applying this method to helical shell models ( i.e. for a class of helical interactions that couple the two helicities in a non separable topology) reveals that a reduction in phase coherence correlates with suppression of intermittency. When inverse cascade scenarios are considered using two different shell models including a non local helical shell model, and a local standard shell model with a modified conservation law, it was shown that a particular phase organization is needed in order to sustain the inverse energy cascade. It was also observed that the PDFs of the triadic phases were peaked in accordance with the basic considerations of the form of the flux, which suggests that a triadic phase of $\pi/2$ and $\pi/2$ maximizes the forward and the inverse energy cascades respectively.
@@ I Introduction
Many natural systems exhibit non-Gaussian fluctuations across a broad range of spatial and temporal scales, implying the presence of extreme fluctuations or rare events, whose dynamical properties and statistical signatures are essential to describe the overall behavior of the system. Examples include seismic activity, financial market crashes and extreme weather events [1, 2, 3]. An archetypical example of such complex, open system is the problem of fully developed turbulence: an out of equilibrium state of fluid or plasma flow, displaying a mixture of hierarchical self organization and chaotic behavior across a large range of scales, maintained in a statistically stationary state by continuous energy injection (e.g. at large scales) and dissipation (e.g. at small scales).
@@
Between forcing and dissipation, turbulent systems can develop an inertial range, spanning many decades, exhibiting self-similar properties heuristically captured by the Richardson cascade picture, and the power law spectrum of Kolmogorov [4].
@@
Since direct numerical simulations (DNS) of turbulent flows over a wide range of scales were historically too costly, various reduced models were developed in the past [5]. These models are capable of reproducing various statistical features of the full system, such as the power law distribution of energy as a function of scale, and the scaling of higher-order structure functions. Among those, shell models [6, 7] stand out in their simplicity and in their ability to mimic certain statistical aspects of different types of turbulent systems including rotating turbulence, passive scalar advection, thermal convection etc. [8, 9, 10]. In the context of plasma physics, shell models have been used to describe magnetohydrodynamic (MHD) turbulence [11, 12], and drift waves in fusion plasmas [13, 14].
@@
Probably the most intriguing – perhaps unexpected – feature of shell models, is their ability to produce the anomalous scaling of the structure functions similar to fully developed turbulence [15, 16, 17]. This suggests that shell models can be a useful tool to capture the statistics of extreme events [18, 19], which are manifested as burst like structures, related to finite time singularities or instantons [20, 21], playing an important role in shaping the underlying intermittency characteristics.
@@
Shell models reduce the complexity of the nonlinear energy transfer governed by triadic interactions drastically, while preserving the key properties of the underlying system such as its scale invariance, the quadratic nature of the nonlinearity and the conservation laws - like Energy and Helicity in the case of three dimensional turbulence. This extreme reduction in the number of degrees of freedom, is a consequence of dividing the Fourier space into a set of geometrically spaced shells and retaining the nonlinear interactions that involve only consecutive shells, thus replacing the turbulent system with a one dimensional chain of connected, self-similar triads.
@@
In order to reintroduce some spatial structure in cascade models, various more sophisticated constructions have been proposed, including nested polyhedra [22], logarithmic discretization (LDM) [23], spiral chains [24] and logarithmic lattices [25, 26, 27]. Most of these models are self-similar by construction, and maintain a constant number of triads across different scales. However, a puzzling feature of these spatially-extended “shell” models is the apparent suppression of intermittency observed in their statistics. Similar loss of intermittency has also been observed in numerical simulations of fractally-decimated 3D turbulence [28] even keeping an overwhelming majority of the modes and dropping only a small percentage.
@@
The effect of a fractal decimation of Fourier modes has been explored also in simpler systems, such as the Burgers equation, where a disruption of the average triadic phase coherence is correlated with a reduction of the intermittency, normally associated with shock formation [29, 30]. This suggests a picture, in which phase alignment between triads enhances energy transfer resulting in strong dissipation events [31, 32]. Extending this analysis to the Navier Stokes equations requires understanding the self-organization of triadic phases in fully developed turbulence, and the definition of a clear order parameter that identifies the interactions involved in the collective behavior responsible for intermittency, which remains a challenge. Part of the difficulty involves dealing with the huge number of degrees of freedom represented by the convolution sum in such a system. One possible way forward may be to reformulate turbulence as a network of Fourier modes [33, 34, 35], which allows us to consider reduced models, such as shell models, as reductions of a complete grid into a network of groups of nodes, in this case represented by shells. Although the individual complex phases of the dynamical variables of a shell model have no physical significance, the fact that the triadic phase coherence along the chain regulates the cascade process, suggests that triadic phase coherence among the underlying Fourier modes, lumped together to form the shells, is related to the resulting shell-to-shell energy transfer.
@@
Thus, to clarify the relationship between intermittency and phase coherence in fully developed turbulence, we investigate this connection in shell models that preserve the same invariants as the Navier-Stokes equations and, in some cases, exhibit a level of intermittency comparable to that of 3D turbulence. Building on our earlier observation that reducing inter-shell spacing in standard shell models increases intermittency considerably [36], we demonstrate a systematic link between triadic phase synchronization (sync) and enhanced intermittency in shell models, consistently with the observations in studies of the Burger’s equation cited above.
@@
Notice that, a shell model can be interpreted as a one dimensional chain of connected triads, each representing a “nonlinear oscillator”. Consecutive triads only differ by a scaling factor related to the inter-shell spacing. This means that by decreasing the inter-shell spacing, (while keeping the conserved quantities unchanged) both the number and the similarity of the oscillators would be increased making it potentially easier to form a phase coherent state. To quantify global phase coherence, and its relation to different levels of anomalous scaling of the structure functions, we employ a global Kuramoto order parameter, which allows us to show that higher intermittency correlates with enhanced phase coherence.
@@
In addition, we observe that phase coherence correlates with the existence of coherent structures in the energy cascade, appearing as localized pulses in shell models. In order to clarify the role of these localized events, we introduce a novel local order parameter –a weighted Kuramoto parameter– which enable us to observe burst-like events propagating across shells in the inertial range and analyze the alignment of triadic phases in forming an “open channel” for energy transfer during these events. We show that extreme energy bursts are consistently preceded by rapid growth in local phase coherence and that more intermittent models display stronger and longer-lasting phase-locked states, thereby facilitating larger energy transfer.
@@
Finally, we consider inverse cascades in shell models, which -apart from particular examples- are usually suppressed by equipartition solutions. We consider these special cases that actually work, in order to study the role of phase coherence and to compare it with the forward cascade scenario. The results reveal that, in the forward and inverse cascade regimes, the PDFs of the triadic phases exhibit peaks at values of $\pi/2$ and $\pi/2$ respectively, which maximize the flux term at fixed amplitude. Notably, in the case of intermittent cascades the peak at $\pi/2$ is much more marked, indicating that coherent events are clearly favored by the system; in contrast, for inverse cascades, the distribution is symmetric. This results are in parallel and complementary with recent results focusing on the relation between phase dynamics and the energy cascade direction [38].
@@
In section II we introduce the shell model framework, focusing on the dynamical equations for phases and amplitudes II.1, discussing role of phase alignments in the energy flux II.2 and presenting the phase statistics for the standard forward intermittent cascade. Next in II.3, we introduce the Kuramoto order parameters to quantify phase coherence in shell models, highlighting how local sync is related to extreme flux events.
@@
In Sec. IV, we extend the analysis to helical shell models, exploring how sync properties vary with shell spacing, focusing on a helical shell model exhibiting suppression of intermittency. Section V is devoted to the role of phase organization in driving inverse cascades. In particular we present an inverse energy cascade for both an elongated helical shell model and a local shell model.
@@
II Shell model as a dynamical system: equations for phases and amplitudes
Motivated by observations of statistical isotropy and scale invariance of the turbulent cascade of fully developed turbulence without obvious sources of anisotropy, the shell model reduction consists in modeling the velocity field as a one-dimensional set of complex variables representing different scales, corresponding to a set of logarithmically spaced wave numbers $k_{n}=k_{0}g^{n}$, where it is customary to take the spacing factor $g=2$.
The evolution equation for the shell variable $u_{n}$, for the GOY (Gledzer-Ohkitani-Yamada) model, which is the standard example for shell models, is usually written as
@@ II Shell model as a dynamical system: equations for phases and amplitudes
where the terms in the square brackets represent the reduction of the convolution sum to a local triadic interaction between three consecutive shells, $d_{n}=\nu k_{n}^{2p}+\mu k_{n}^{-2q}$ is the general form of (hyper/hypo)-viscosity, and $f_{n}$ denotes external forcing, localized at a shell $n_{f}$, or a small set of consecutive shells around it, corresponding to the typical energy injection scale $1/k_{n_{f}}$.
Here the coefficients $a,b,c$ are chosen in such a way that they preserve typical inviscid invariants of the underlying physical system. In the case of 3D Navier-Stokes turbulence, it is the energy and the helicity that are conserved, which are defined for the shell models respectively as
@@
Enforcing the conservation of these two quantities one obtains, $a=g+1$, $b=g^{-1}-g$ and c=1gc=-1-g, as a function of the inter-shell spacing $g$. It is also common to factor out $ak_{n+1}$ in the nonlinear term of Eq. (1) so that the other two coefficients corresponding to $b$ and $c$ can be written as $\left(g^{-2}-g^{-1}\right)$ and $g^{-3}$ respectively. The common choice of $g=2$ would then give 1/4-1/4 and 1/8-1/8 for the second and third coefficients respectively. Another common choice in the literature is $g=\varphi$ ( the golden ratio), which gains significance, from the derivation of shell models based on a decimation of Fourier space, which requires the existence of triadic interations between the characteristic wave numbers of the shells (specifically solution to $\boldsymbol{k}_{n-1}+\boldsymbol{k}_{n}+\boldsymbol{k}_{n+1}=0$, which imposes the geometrical constrain $g\leq\varphi$ ). While there are some examples of studies of shell models varying their nonlinear coupling coefficients [39, 40], most of these studies are carried out in such a way that varying the coefficients change the conservation law, and hence the effective dimensionality of the underlying physical model, instead of keeping the conservation laws fixed and varying the logarithmic spacing for a fixed model. Exploring different shell spacings is also important for making the connection to a perspective based on Fourier space decimation, such as log-lattices [26], LDM [23], or helical shell models with different types of elongations [41].
@@
Note that a standard shell model with $N$ nodes has N2N-2 interacting triads. Since the equation for a single unconnected triad is equivalent to that of a spinning top, giving rise to nonlinear oscillations that can be represented in terms of Jacobi elliptic functions, each triad can be considered as a nonlinear oscillator. Furthermore since the shell model consists of self-similar triads that are scaled by the same factor, the consecutive nonlinear oscillators that are coupled are also self-similar, maintaining the same scaling factor. The energy flux through such a chain of triads then requires some sort of synchronization of the phases of those consecutive triads. Decreasing the shell spacing is thus key to understanding the synchronization behavior of shell models. Since it simultaneously allows us to increase $N$, considering the large $N$ limit, important for synchronization, and reduce the ratio of the coefficients of two consecutive triads allowing them to have closer natural nonlinear frequencies. Note that the usual GOY model with $g=2$, covers roughly seven decades in $k$-space with $N=24$, and it is not feasible to increase that number substantially without decreasing $g$, since the power law solution would start going under machine precision levels.
@@
Thus, varying the $g$ parameter, while increasing $N$ in order to cover the same wave-number range, allows us to modify the affinity of the triadic network of the shell model to synchronization (i.e. synchronizability). Since the intermittency in shell models is a consequence of its time evolution, one would expect this to have significant impact on intermittency as well. Hence, we present below, a detailed study introducing specific methods for observing various aspects of local and global synchronization, and the implications of these observations for the turbulent cascade as well as intermittency, mainly by varying the $g$ parameter to control its sychronizability.
II.1 Phase Dynamics in Shell Models
@@
which leaves the triadic phases invariant for any constant $\delta.$ The consequence of this $U(1)$ symmetry is that single shell phases $\theta_{n}$ are not uniquely determined by the triadic phases, and as also numerically confirmed, they remain uniformly distributed (i.e. “random”). Thus, the appropriate quantities, whose dynamics and statistics are physically relevant are the triadic phases $\varphi_{n}$, rather than single shell phases $\theta_{n}$.
@@
$+a\Lambda_{n+2}^{n+1}\cos\varphi_{n+2}+b\Lambda_{n+1}^{n+1}\cos\varphi_{n+1}+c\Lambda_{n}^{n+1}\cos\varphi_{n}$
@@
$\Lambda_{\ell}^{n}\equiv k_{\ell}\frac{U_{\ell-1}U_{\ell}U_{\ell+1}}{U_{n}^{2}}.$
Note that Eq. (5) involves the triadic phases of five consecutive triads from n2n-2 to $n+2$, which in turn include contributions from seven consecutive shells from $n+3$ to n3n-3, making this coupling longer range compared to the shell amplitude equation, which involves only the neighboring triadic phases (i.e. $\varphi_{j}$ with j=n1,n,n1j=n-1,n,n-1) and contributions from shells from n2n-2 to $n+2$. Note that these couplings can also be interpreted (and generalized to more complex interactions) within the network representation of the shell models, see e.g. Ref. 34, where the evolution of a given triad can be computed as a sum over all the triads connected through shared nodes in the network topology.
II.2 The Role of Triadic Phase on Energy Flux
The phase dynamics of a shell model with $N$ shells can therefore be considered as being equivalent to that of a chain of N2N-2 nonlinear oscillators, each representing a triad, where the coupling terms are determined by a combination of the $N$ self-consistently evolving amplitudes and the interaction coefficients. When forcing is applied near the first few shells and dissipation occurs at high-$n$, the chain of oscillators is constrained to accommodate the energy flux from low to high $n$. In particular if the phases are initially distributed in such a way that there is no initial net flux, the amplitudes will grow until the phases will reorganize themselves in order to accommodate the spectral flux implied by forcing and dissipation, basically because of the free energy associated with such a non-uniform energy distribution. This means that the energy injection at low-$n$ drives the phases to dynamically self-organize in order to sustain the cascade mechanism, possibly reaching an eventual steady state.
In order to illustrate this steady state, let us consider the form of the flux in a shell model. The energy cascades to small scales due to nonlinear interactions, and the energy transfer in a given triad centered at shell $n$ is defined as a local spectral energy flux, expressed in terms of the three shell correlation $\chi_{n}=U_{n-1}U_{n}U_{n+1}e^{-i\varphi_{n}}$ as:
$\begin{array}[]{cc}\Pi_{n}^{E}=&2\text{Im}\left(ak_{n+1}\chi_{n+1}-ck_{n}\chi_{n}\right).\end{array}$ (6)